# and quantum information science

### v2

@@ 7.4 Generalizations of the holographic proposals
Next, let us consider the switchback effect. In the case of a finite-energy shock wave perturbing a black hole solution in asymptotically de Sitter spacetime, the standard complexity conjectures (CV, CV2.0 and CA) admit a plateau around $t=0$ when complexity is approximately constant [359, 360], similar to the AdS case. [^36] The de Sitter shock waves considered in [359, 360] correspond to the (spherically symmetric) ejection of mass into the cosmological horizon. As a consequence, the cosmological horizon is pushed further away from an observer sitting at some fixed radius. While this setting may seem somewhat peculiar, such geometries are consistent with the null energy condition and provide a simple setup in which complexity can be studied. Notably, the duration of this regime increases when the shock crosses the stretched horizon at earlier times (see figs. 23, 25 in [359]), and corresponds to special geometrical configurations that arise because the Penrose diagram of de Sitter spacetime grows taller when a null pulse carrying positive energy is inserted in the bulk [361]. The critical time at which hyperfast growth occurs in CV, CV2.0 and CA conjectures is always delayed by the presence of a shock wave moving along the cosmological horizon [362] (see also [363]). Reference [364] showed that the codimension-one CAny observables displaying linear growth at late times present a delay in this evolution satisfying the same structure obtained in Eq. (123) in the AdS case. All these behaviors support the existence of a switchback effect for the holographic conjectures applied to de Sitter spacetime. It would be useful to match the above behaviors with those of appropriate circuit models, similarly to the arguments of section 4.1.
@@ 7.5 Quantitative matches relating spread complexity and gravitational observables
In this section, we will review proposals for quantitative matches between notions of complexity in the boundary theory and geometric observables in the dual bulk. The basic challenge is to frame the definition of complexity in terms of the holographic dictionary [365, 366] in order to get a corresponding quantity in the bulk gravity. A notable recent success is the matching between the length of wormholes in Jackiw-Teitelboim gravity and spread complexity in a dual Double Scaled SYK model.
@@ 7.5.1 Spread complexity in double-scaled SYK model equals wormhole length
where $\lambda=\frac{2p^{2}}{N}$. The scaling of the couplings is tuned so that the density of states is bounded between $-2\mathcal{J}$ and $2\mathcal{J}$ in the large-$N$ limit with $p$ fixed (see [181] for a review). Here we focus on the so-called Double Scaled SYK (DSSYK) model in which $N,p\to\infty$ with $\lambda=2p^{2}/N$ fixed (see the review [367]). Below, we will also refer to a triple scaling limit that is defined by $\lambda\ll 1$ and energies $E/\mathcal{J}\ll 1$. In this limit the theory is governed by the Schwarzian action [180, 33].
@@
where we introduced the notation $\psi_{I}=\psi_{i_{1}}\psi_{i_{2}}\cdots\psi_{i_{p}}$. We have only written the even moments because the odd moments vanish. To finish the computation we have to Wick contract the multi-indexed fermion terms in pairs. As discussed in [368, 32], the Feynman diagrams describing these contractions are called chord diagrams and can be visualized by placing $2k$ points on the circumference of a circle, and then contracting the dots in pairs. Carrying out these contractions, we arrive at the moments
@@
where $q=e^{-\lambda}$ The second equality expresses the moments in terms of a tridiagonal transfer matrix $T$ acting on $\lvert 0\rangle$, an auxiliary state with zero chords, that represents the thermofield double state at infinite temperature [32]. Essentially, by construction $T$ is the tridiagonalized Hamiltonian; the apparently auxiliary chord basis is precisely the Krylov basis of Sec. 6. Now, following the moment method, we can immediately read off the Lanczos coefficients. They are
@@
We can then solve the Schrodinger equation, and compute spread complexity. In the double scaling limit $N\to\infty$, and hence the Hilbert space is infinite-dimensional. Correspondingly, using the Lanczos coefficients in (132), we find a spread complexity that grows linearly forever (see Sec. 6). Meanwhile, we can apply an additional low-energy scaling limit [32]. This so-called triple scaling limit is effectively a continuum limit for the chord basis in which $T$, the tridiagonalized Hamiltonian, reduces to a boundary description of a bulk Jackiw-Teitelboim (JT) gravity. In this limit, the authors of [33, 34] showed that the quantity calculating the length of the wormhole in the TFD state of JT gravity is precisely the spread complexity discussed above. This construction establishes a precise correspondence between wormhole length in gravity and spread complexity in this limit of JT gravity and the dual DSSYK model.
However, this construction leads to a puzzle. At any finite $N$, the Hilbert space should be finite-dimensional, and hence the spread complexity should saturate at late times, together with the quantum analog of wormhole length. How can we see this? Fortunately, the results of [368] demonstrate that the moments of the SYK Hamiltonian described above arise from the distribution
@@
In other words, this is the density of states at a given $q$. Following the discussion in Sec. 6.1.4, we can directly compute the Lanczos spectrum from this density of states at large but finite $N$ via an integral formula. This procedure was carried out in [249, 35], where the authors demonstrated that the Lanczos coefficients for the SYK model match the computations of [32] (described above) for an initial range of Krylov indices, but then decline systematically to zero. The decline occurs over the whole range of Krylov indices up to the dimension of the Hilbert space. with a slope of $1/N$, where $N$ is the dimension of the Hilbert space. This descent to zero for the Lanczos coefficients causes the spread complexity to saturate at late times. If we maintain the identification between wormhole length and spread complexity that applies at early times, we would say that the wormhole length is saturating at late times. From the bulk JT gravity point of view, this decline thus corresponds to a non-perturbative quantum effect in gravity that corrects the steady increase in the Lanczos coefficients and wormhole size expected from the early-time classical description of the system. Of course, at late times we could imagine alternative non-perturbative definitions of the wormhole length in JT gravity that also agree with the early-time identification with spread complexity. Interesting recent developments along this line are also presented in [31, 369, 269] which were discussed in Sec. 7.2.
@@
In a related vein, Ref. [370] utilized new developments identifying the gravity dual of DSSYK [371] to show that the relation between spread complexity and the bulk volume extends past the limits used in the studies of [34]. In particular, this relation extends to the full DSSYK regime at arbitrary temperatures. The key finding of [370] was that the match with the volume on the gravity side required computing the spread complexity for the preparation of a finite-temperature state using the Euclidean path integral. This is opposed to including the Euclidean state preparation in the reference state, and then assigning non-vanishing complexity only to the purely Lorentzian evolution. Furthermore, since the calculations in [370] also applied to the quantum regime, the authors were able to extract the form of the leading correction to the CV proposal originating from the contributions of quantum fields in a two-dimensional bulk geometry. This correction was crucially needed for the quantitative match with the boundary spread complexity. Interestingly, the correction exhibits a behavior compatible with the expectations for the spread complexity of bulk quantum fields, but a precise relation remains unknown.
@@
These findings suggest that Krylov-basis definitions of complexity are natural from a bulk perspective, and could provide a non-perturbative definition of bulk volumes. A crucial next step would be to extend these studies to higher dimensions, beginning with AdS3. While the powerful exact results available in DSSYK may not extend directly, recent progress with Virasoro TQFTs [372] could offer insights into the bulk Hilbert space and its relation to the Krylov basis. Another approach could be to exploit the observation in [373] that at late times the volume of an extremal slice through the ERB of any spherically-symmetric black hole reduces to a geodesic length in an effective two-dimensional JT gravity theory. This suggests that we may be able to directly export the results relating spread complexity and wormholes in JT gravity to higher dimensions.
@@ 7.5.2 Rate of spread-complexity growth and proper momentum
The above comparisons between the Krylov basis definitions of complexity and holography were made in the context of low-dimensional toy models. This progress was possible due to the explicit definition of the thermofield-double state in the double-scaled SYK model using a transfer matrix and its $q$–deformed algebra. However, extending this approach to higher-dimensional holographic setups may be challenging. Nevertheless, several universal insights about operator growth and spread complexity can already be tested in higher dimensions. One of the most concrete examples is the relation between the growth rate of operator complexity (or size) and the radial momentum of its dual massive particle in AdS spacetime. Intuitively, we expect that the time derivative of complexity (the rate of complexity) of the Heisenberg operator $\mathcal{O}(t)$ should be proportional to the radial momentum of the particle
@@
Since $\mathcal{C}$ is the mean position in the Krylov chain, this relation is telling us that the rate of progress along this chain is dual to the radial momentum. This sort of relation has been argued for in various ways in [49, 374, 375, 376, 248], including qualitative models of operator growth using epidemic models [377]. Furthermore, Refs. [378, 379, 380] essentially proved this relation under the assumption of the CV proposal. However, at that time, a precise definition of complexity satisfying Eq. (134) was missing.
By now, there are setups where this formula can be precisely evaluated on both sides of the holographic correspondence. Firstly, in the SYK model at large $q$, one can directly compute the expectation value of the size operator [165, 381] that can be defined in terms of two copies of fermions (left and right in the thermofield-double construction, denoted with subscripts $l$ and $r$ respectively) as
@@
and in [381], authors showed that its rate of change in time (i.e., time derivative) is proportional to the radial momentum of a particle in AdS2 (see also [382]).
Recently, this correspondence was also confirmed between the spread complexity in 2D CFTs and the proper radial momentum of particles in AdS3 spacetime. Reference [383] studied a local operator quench where a CFT state is locally excited by a primary operator with conformal dimension $\Delta$, and evolved unitarily with the Hamiltonian. Holographically, this setting corresponds to a particle with mass $m=\Delta$ propagating from the boundary towards the bulk, with the particle’s initial position corresponding to the operator’s energy regulator $\varepsilon$. Interestingly, the rate of spread complexity for this dynamical state is related to the particle’s momentum via
@@
However, the key observation is that the momentum must be computed in the proper radial distance coordinate $\rho$ (for example, this relation would not hold in Poincaré coordinates). This relation between Krylov complexity and momentum has also been analyzed from a slightly different perspective in [384, 385], and the connection between the position on the Krylov chain $n$ and proper radial distance in the bulk was further substantiated in [386].
Finally, the evolution of spread complexity following local quench protocols in 2D CFTs was studied in [155]. The initial state is given by two thermal states on semi-infinite lines joined together at $x=0$ and evolved with the CFT Hamiltonian. The crucial object, as discussed in Sec. 6, is the return amplitude that reads
@@
These are examples of the $\mathbb{R}$) analytical solutions discussed in Sec. 6.3.1. We can see that, similarly to the SYK example, the large–$c$ limit effectively constrains the range of $n$. Indeed, if we first extract the naive large–$c$ limit, Lanczos coefficients resemble the Heisenberg-Weyl example with $b_{n}\sim\sqrt{n}$ and quadratic growth of complexity. This is correct as long as $n\ll c/8$. If $n$ is comparable or larger than $c/8$, we have to take the full answer first.
@@
This local quench setup is described holographically using the AdS/BCFT correspondence [387], where the end-of-the-world (EOW) brane is time-dependent. Interestingly, the tip of the EOW brane, which probes the deepest part of the AdS bulk, follows a geodesic of a massive particle (with heavy mass $m=c/16$). The radial momentum of this particle, computed in the proper distance coordinate, precisely satisfies
@@
These explicit checks suggest that, depending on the initial state, it may be possible to engineer precise holographic setups where the spread complexity rate of change can be directly matched with a quantity in gravity and then integrated. The overall dependence on the central charge in (139) is particularly encouraging and enters similarly as in the single-interval entanglement entropy in 2D CFTs that can be reproduced from the Ryu-Takayanagi formula [5].
To finish, let us point out that an intriguing perspective on the Krylov basis was also presented in [156], where, using Wigner functions, the authors argued that the Krylov basis is ”the most classical” basis in Quantum Field Theory, potentially making it a natural candidate for describing the evolution of semi-classical states in the bulk code subspace.

### v3

@@ 7.4 Generalizations of the holographic proposals
Next, let us consider the switchback effect. In the case of a finite-energy shock wave perturbing a black hole solution in asymptotically de Sitter spacetime, the standard complexity conjectures (CV, CV2.0 and CA) admit a plateau around $t=0$ when complexity is approximately constant Baiguera:2023tpt, Baiguera:2024xju, similar to the AdS case. [^39] The de Sitter shock waves considered in Baiguera:2023tpt, Baiguera:2024xju correspond to the (spherically symmetric) ejection of mass into the cosmological horizon. As a consequence, the cosmological horizon is pushed further away from an observer sitting at some fixed radius. While this setting may seem somewhat peculiar, such geometries are consistent with the null energy condition and provide a simple setup in which complexity can be studied. Notably, the duration of this regime increases when the shock crosses the stretched horizon at earlier times (see figs. 23, 25 in Baiguera:2023tpt), and corresponds to special geometrical configurations that arise because the Penrose diagram of de Sitter spacetime grows taller when a null pulse carrying positive energy is inserted in the bulk Gao:2000ga. The critical time at which complexity starts growing significantly in CV, CV2.0 and CA conjectures is always delayed by the presence of a shock wave moving along the cosmological horizon Anegawa:2023dad (see also Faruk:2025bed). More precisely, the duration of the above-mentioned plateau region is $t_{pl}=4(t_{w}-t_{*})$ where $t_{w}$ is the time at which the shockwave is inserted and $t_{*}\sim\frac{1}{T_{C}}\log S_{C}$ is the analogue of the scrambling time (with $T_{C},S_{C}$ the temperature and entropy of the cosmic horizon for the Schwarzschild-de Sitter geometry to which the shock is inserted) Baiguera:2023tpt, Baiguera:2024xju. This result is similar to the one found in the AdS case. Reference Aguilar-Gutierrez:2023pnn further showed that the codimension-one CAny observables displaying linear growth at late times present a delay in this evolution satisfying the same structure obtained in Eq. (123) in the AdS case. All these behaviors support the existence of a switchback effect for the holographic conjectures applied to de Sitter spacetime. It would be useful to match the above behaviors with those of appropriate circuit models, similarly to the arguments of section 4.1. In this regard, it is important to note that the nature of scrambling in de Sitter spacetime has been a subject of debate, see e.g., Refs. Aalsma:2020aib, Geng:2020kxh, Chapman:2021eyy, Milekhin:2024vbb.
@@ 7.5 Quantitative matches relating spread complexity and gravitational observables
In this section, we will review proposals for quantitative matches between notions of complexity in the boundary theory and geometric observables in the dual bulk. The basic challenge is to frame the definition of complexity in terms of the holographic dictionary Gubser:1998bc, Witten:1998qj in order to get a corresponding quantity in the bulk gravity. A notable recent success is the matching between the length of wormholes in Jackiw-Teitelboim gravity and spread complexity in a dual Double Scaled SYK model.
@@ 7.5.1 Spread complexity in double-scaled SYK model equals wormhole length
where $\lambda=\frac{2p^{2}}{N}$. The scaling of the couplings is tuned so that the density of states is bounded between $2\mathcal{J}$ and $2\mathcal{J}$ in the large-$N$ limit with $p$ fixed (see Sarosi:2017ykf for a review). Here we focus on the so-called Double Scaled SYK (DSSYK) model in which $N,p\to\infty$ with $\lambda=2p^{2}/N$ fixed (see the review Berkooz:2024lgq). Below, we will also refer to a triple scaling limit that is defined by $\lambda\ll 1$ and energies $E/\mathcal{J}\ll 1$. In this limit the theory is governed by the Schwarzian action Maldacena:2016hyu, Lin:2022rbf.
@@
where we introduced the notation $\psi_{I}=\psi_{i_{1}}\psi_{i_{2}}\cdots\psi_{i_{p}}$. We have only written the even moments because the odd moments vanish. To finish the computation we have to Wick contract the multi-indexed fermion terms in pairs. As discussed in Erdos:2014zgc, Berkooz:2018qkz, the Feynman diagrams describing these contractions are called chord diagrams and can be visualized by placing $2k$ points on the circumference of a circle, and then contracting the dots in pairs. Carrying out these contractions, we arrive at the moments
@@
where $q=e^{-\lambda}$. The second equality expresses the moments in terms of a tridiagonal transfer matrix $T$ acting on $\lvert 0\rangle$, an auxiliary state with zero chords, that represents the thermofield double state at infinite temperature Berkooz:2018qkz. Essentially, by construction $T$ is the tridiagonalized Hamiltonian; the apparently auxiliary chord basis is precisely the Krylov basis of Sec. 6. Now, following the moment method, we can immediately read off the Lanczos coefficients. They are
@@
We can then solve the Schrodinger equation, and compute spread complexity. In the double scaling limit $N\to\infty$, and hence the Hilbert space is infinite-dimensional. Correspondingly, using the Lanczos coefficients in (132), we find a spread complexity that grows linearly forever (see Sec. 6). Meanwhile, we can apply an additional low-energy scaling limit Berkooz:2018qkz. This so-called triple scaling limit is effectively a continuum limit for the chord basis in which $T$, the tridiagonalized Hamiltonian, reduces to a boundary description of a bulk Jackiw-Teitelboim (JT) gravity. In this limit, the authors of Lin:2022rbf, Rabinovici:2023yex showed that the quantity calculating the length of the wormhole in the TFD state of JT gravity is precisely the spread complexity discussed above. This construction establishes a precise correspondence between wormhole length in gravity and spread complexity in this limit of JT gravity and the dual DSSYK model.
However, this construction leads to a puzzle. At any finite $N$, the Hilbert space should be finite-dimensional, and hence the spread complexity should saturate at late times, together with the quantum analog of wormhole length. How can we see this? Fortunately, the results of Erdos:2014zgc demonstrate that the moments of the SYK Hamiltonian described above arise from the distribution
@@
In other words, this is the density of states at a given $q$. Following the discussion in Sec. 6.1.4, we can directly compute the Lanczos spectrum from this density of states at large but finite $N$ via an integral formula. This procedure was carried out in Nandy:2024zcd, Balasubramanian:2024lqk, where the authors demonstrated that the Lanczos coefficients for the SYK model match the computations of Berkooz:2018qkz (described above) for an initial range of Krylov indices, but then decline systematically to zero. The decline occurs over the whole range of Krylov indices up to the dimension of the Hilbert space. This descent to zero for the Lanczos coefficients causes the spread complexity to saturate at late times. If we maintain the identification between wormhole length and spread complexity that applies at early times, we would say that the wormhole length is saturating at late times. From the bulk JT gravity point of view, this decline thus corresponds to a non-perturbative quantum effect in gravity that corrects the steady increase in the Lanczos coefficients and wormhole size expected from the early-time classical description of the system. Of course, at late times we could imagine alternative non-perturbative definitions of the wormhole length in JT gravity that also agree with the early-time identification with spread complexity. Interesting recent developments along this line are also presented in Iliesiu:2021ari, Iliesiu:2024cnh, Miyaji:2025yvm which were discussed in Sec. 7.2.
@@
In a related vein, Ref. Heller:2024ldz utilized new developments identifying the gravity dual of DSSYK Blommaert:2024ydx to show that the relation between spread complexity and the bulk volume extends past the limits used in the studies of Rabinovici:2023yex. In particular, this relation extends to the full DSSYK regime at arbitrary temperatures. The key finding of Heller:2024ldz was that the match with the volume on the gravity side required computing the spread complexity for the preparation of a finite-temperature state using the Euclidean path integral. This is opposed to including the Euclidean state preparation in the reference state, and then assigning non-vanishing complexity only to the purely Lorentzian evolution. Furthermore, since the calculations in Heller:2024ldz also applied to the quantum regime, the authors were able to extract the form of the leading correction to the CV proposal originating from the contributions of quantum fields in a two-dimensional bulk geometry. This correction was crucially needed for the quantitative match with the boundary spread complexity. Interestingly, the correction exhibits a behavior compatible with the expectations for the spread complexity of bulk quantum fields, but a precise relation remains unknown.
@@
These findings suggest that Krylov-basis definitions of complexity are natural from a bulk perspective, and could provide a non-perturbative definition of bulk volumes. A crucial next step would be to extend these studies to higher dimensions, beginning with AdS3. While the powerful exact results available in DSSYK may not extend directly, recent progress with Virasoro TQFTs Collier:2023fwi could offer insights into the bulk Hilbert space and its relation to the Krylov basis. Another approach could be to exploit the observation in Gautason:2025ryg that at late times the volume of an extremal slice through the ERB of any spherically-symmetric black hole reduces to a geodesic length in an effective two-dimensional JT gravity theory. This suggests that we may be able to directly export the results relating spread complexity and wormholes in JT gravity to higher dimensions.
@@ 7.5.2 Rate of spread-complexity growth and proper momentum
The above comparisons between the Krylov basis definitions of complexity and holography were made in the context of low-dimensional toy models. This progress was possible due to the explicit definition of the thermofield-double state in the double-scaled SYK model using a transfer matrix and its qq–deformed algebra. However, extending this approach to higher-dimensional holographic setups may be challenging. Nevertheless, several universal insights about operator growth and spread complexity can already be tested in higher dimensions. One of the most concrete examples is the relation between the growth rate of operator complexity (or size) and the radial momentum of its dual massive particle in AdS spacetime. Intuitively, we expect that the time derivative of complexity (the rate of complexity) of the Heisenberg operator $\mathcal{O}(t)$ should be proportional to the radial momentum of the particle
@@
Since $\mathcal{C}$ is the mean position in the Krylov chain, this relation is telling us that the rate of progress along this chain is dual to the radial momentum. This sort of relation has been argued for in various ways in Susskind:2014jwa, Susskind:2018tei, Susskind:2019ddc, Magan:2018nmu, Lin:2019kpf, including qualitative models of operator growth using epidemic models Susskind:2020gnl. Furthermore, Refs. Barbon:2020uux, Barbon:2020olv, Barbon:2019tuq essentially proved this relation under the assumption of the CV proposal. However, at that time, a precise definition of complexity satisfying Eq. (134) was missing.
By now, there are setups where this formula can be precisely evaluated on both sides of the holographic correspondence. Firstly, in the SYK model at large $q$, one can directly compute the expectation value of the size operator Roberts:2018mnp, Lin:2019qwu that can be defined in terms of two copies of fermions (left and right in the thermofield-double construction, denoted with subscripts $l$ and $r$ respectively) as
@@
and in Lin:2019qwu, authors showed that its rate of change in time (i.e., time derivative) is proportional to the radial momentum of a particle in AdS2 (see also Brown:2018kvn).
Recently, this correspondence was also confirmed between the spread complexity in 2D CFTs and the proper radial momentum of particles in AdS3 spacetime. Reference Caputa:2024sux studied a local operator quench where a CFT state is locally excited by a primary operator with conformal dimension $\Delta$, and evolved unitarily with the Hamiltonian. Holographically, this setting corresponds to a particle with mass $m=\Delta$ propagating from the boundary towards the bulk, with the particle’s initial position corresponding to the operator’s energy regulator $\varepsilon$. Interestingly, the rate of spread complexity for this dynamical state is related to the particle’s momentum via
@@
However, the key observation is that the momentum must be computed in the proper radial distance coordinate $\rho$ (for example, this relation would not hold in Poincaré coordinates). This relation between Krylov complexity and momentum has also been analyzed from a slightly different perspective in Fan:2024iop, He:2024pox, and the connection between the position on the Krylov chain $n$ and proper radial distance in the bulk was further substantiated in Dodelson:2025rng.
Finally, the evolution of spread complexity following local quench protocols in 2D CFTs was studied in Caputa:2025dep. The initial state is given by two thermal states on semi-infinite lines joined together at $x=0$ and evolved with the CFT Hamiltonian. The crucial object, as discussed in Sec. 6, is the return amplitude that reads
@@
[…]
@@
This local quench setup is described holographically using the AdS/BCFT correspondence Takayanagi:2011zk, where the end-of-the-world (EOW) brane is time-dependent. Interestingly, the tip of the EOW brane, which probes the deepest part of the AdS bulk, follows a geodesic of a massive particle (with heavy mass $m=c/16$). The radial momentum of this particle, computed in the proper distance coordinate, precisely satisfies
@@
These explicit checks suggest that, depending on the initial state, it may be possible to engineer precise holographic setups where the spread complexity rate of change can be directly matched with a quantity in gravity and then integrated. The overall dependence on the central charge in (139) is particularly encouraging and enters similarly as in the single-interval entanglement entropy in 2D CFTs that can be reproduced from the Ryu-Takayanagi formula Ryu:2006bv.
To finish, let us point out that an intriguing perspective on the Krylov basis was also presented in Basu:2024tgg, where, using Wigner functions, the authors argued that the Krylov basis is ”the most classical” basis in Quantum Field Theory, potentially making it a natural candidate for describing the evolution of semi-classical states in the bulk code subspace.